OpenStudy (anonymous):

Calculate the maximum energy of an electron in a cathode ray tube operated at a plate voltage of 50 V. Express your answer in units of Joules.

OpenStudy (aaronq):

you can use these equations to find the KE of the electrons emitted. \(\lambda=\dfrac{hc}{eV}\) \(\lambda=\dfrac{h}{mv}\) \(KE=\dfrac{1}{2}mv^2\)